# Dynamic spillovers and investment strategies across artificial intelligence ETFs, artificial intelligence tokens, and green markets

### v1

@@ 5.2 Multivariate portfolio analysis
Table 3 show the results for MVP, MCP and MCoP. In the MVP strategy, the relatively high HE values (IRBO: 0.905, ROBT: 0.890) indicate that multivariate portfolios are effective in reducing the risk of AI ETFs. Even though AI ETFs are the primary transmitters of shocks in the return transmission analysis, their weights remain relatively low (IRBO: 2.4%, ROBT: 3.8%). This suggests that their role in the MVP portfolio is limited. In the MCP, the allocation to AI ETFs increases significantly (IRBO: 6.0%, ROBT: 12.5%), indicating an enhanced contribution to the portfolio. However, their HE values turn negative (IRBO: $-2.215$, ROBT: $-2.728$), indicating that this portfolio fails to effectively mitigate volatility in AI ETFs and may even increase risk. In the MCoP method, the allocation to AI ETFs further increases (IRBO: 10.5%, ROBT: 11.5%), reinforcing their role within the portfolio. Nevertheless, their HE values continue to decline (IRBO: $-3.375$, ROBT: $-4.073$). Overall, despite Figs. 2, 4 and 6 indicating that AI ETFs are major risk spillovers, their weights in these portfolios remain relatively limited.
@@
Next, we analyze the role of AI Tokens across different portfolio strategies. In the MVP, AI tokens (FET, FIL, ICP, NEAR, RENDER) have minimal weights, each below 1%, limiting their overall impact on the portfolio. Furthermore, their HE values exceed 0.99, suggesting that this investment strategy effectively minimizes their risk. In the MCP portfolio, despite a decrease in HE values (ranging between 0.724 and 0.854), the weights of AI Tokens increase significantly (e.g., RENDER: 12.6%, FET: 11.6%), indicating their enhanced role within the portfolio. In the MCoP, the weights of AI Tokens further increase, reflecting their rising importance in this investment strategy. Their HE values remain above 0.7 (ranging between 0.705 and 0.802), suggesting that this strategy contribute to volatility reduction of AI tokens, though to a lesser extent than in the MVP. Overall, AI Tokens exhibit higher allocation ratios for both MCP and MCoP, making them core assets, whereas their allocation in the MVP remains low due to higher volatility, limiting their overall influence.
@@
Finally, we examine the performance of green assets in the portfolios. Green assets hold the highest allocation in the MVP, particularly green bonds (SPGB: 91.2%, SPGTCED: 0.7%). Green assets remain the core holdings in the MVP, with the highest allocation. SPGB holds an overwhelming weight of 91.2%, while SPGTCED accounts for 0.7%, likely reflecting their respective roles as a risk receiver and a risk contributor. Accordingly, SPGTCED has an HE value of 0.884, demonstrating a significantly stronger risk reduction effect compared to SPGB, which has a negative HE value of --0.05. From the MCP procedure, the weight of green assets declines significantly (SPGB: 19.7%, SPGTCED: 13.2%, total: 32.9%), and their HE values continue to decline, indicating a weakening risk reduction in this portfolio. In the MCoP, the allocation to green assets decreases further (SPGB: 11.2%, SPGTCED: 11.0%, total: 22.2%), with HE values continuing to drop, suggesting a further reduction in their role. Despite Figures 2, 4 and 6 demonstrating that clean energy transmits shocks to other markets, its weight in the portfolio is not dominant. This finding aligns with previous literature [15], which also found that while clean energy is a risk contributor, its allocation in investment portfolios is not necessarily the highest, whereas green bonds tend to account for a larger proportion of portfolio allocation.
@@
SPGB 0.197 0.034 0.139 0.261 --34.611 0.000
SPGTCED 0.132 0.034 0.061 0.176 --2.927 0.000
MCoP IRBO 0.105 0.112 0.000 0.250 --3.375 0.000
ROBT 0.115 0.120 0.000 0.250 --4.073 0.000
@@
SPGB 0.112 0.118 0.000 0.250 --47.457 0.000
SPGTCED 0.110 0.113 0.000 0.250 --4.343 0.000
Interestingly, in the case of the MVP, asset allocations, except for AI tokens, are directly opposite to their roles. Risk receivers have relatively higher weights, while risk transmitters have lower weights. The results of MCP and MCoP are quite similar, possibly because both focus on reducing correlations across assets. Both portfolios reduce the risk of AI tokens but increase the volatility of the other two asset classes.
5.3 Portfolio performance
We illustrate the cumulative portfolio returns in Fig. 8. This figure presents the cumulative returns of three portfolio strategies: MVP (black line), MCP (red line), and MCoP (light blue line). MVP remains relatively stable with minimal fluctuations, indicating a low-risk but low-return profile, while MCP and MCoP show significant growth, especially from early 2023 onward. In the early period, all portfolios experience negative returns, but MCP and MCoP recover strongly and reach cumulative returns above 1.0 by early 2025, whereas MVP remains around the zero level. MCP and MCoP follow similar trends, with MCoP generally achieving slightly higher peaks, suggesting better performance in specific market conditions. Other studies on investment strategy comparisons also documents negative cumulative portfolio returns [27, 28].
Figure 8: Cumulative returns of different portfolios.
Finally, the Sharpe ratios of the three portfolios are shown in Table 4. The MCP emerges as the most optimal strategy, as it achieves the highest return and Sharpe ratios across all evaluated metrics. While the MVP effectively minimizes volatility, its negative return and Sharpe ratios render it an unattractive investment option. The MCoP offers a moderate risk-return profile, yet it fails to surpass MCP. Therefore, for investors seeking an optimal balance between risk and return, the MCP portfolio represents the most advantageous selection. This finding is similar with Ziadat et al. [29], who also observe a negative Sharpe ratio for the MVP, while the MCP outperforms both the MVP and MCoP.
@@ 5.3 Portfolio performance
           MVP            MCP            MCoP
           Return            --0.0165            0.2664            0.2438
           StdDev            0.0835            0.4860            0.5669
           Sharpe Ratio (StdDev)            --0.1978            0.5482            0.4300
           Sharpe Ratio (VaR)            --1.9865            5.8477            4.8230
           Sharpe Ratio (CVaR)            --1.4889            4.2027            3.8360
@@ 6 Conclusion
This study employs the $R^{2}$ decomposition to examine the risk spillover dynamics among AI ETFs, AI tokens, and green assets over the period from May 25, 2021, to January 31, 2025. Our findings reveal that contemporaneous spillovers dominate the total connectedness index, while lagged spillovers remain relatively stable. Furthermore, AI ETFs and clean energy emerge as risk transmitters, whereas AI tokens and green bonds primarily act as risk receivers. In addition, AI tokens are challenging to hedge and offer little benefit in hedging other assets when considering bilateral hedge ratios. However, clean energy and AI ETFs exhibit stronger hedging effectiveness regardless of the long or short position taken. Additionally, we analyze the performance of three portfolio strategies: MVP, MCP, and MCoP. The results show that all the approaches reduce volatility of AI tokens significantly. However, only the MVP lowers the risk of other assets, while the other two strategies actually increase their risk. Finally, MCP exhibits superior performance than the other portfolios. In this investment strategy, green bonds have the largest allocation, while IRBO has the smallest.
@@
Our work has the following implications: First, Investors should consider the high spillover risks associated with AI ETFs and clean energy. Although the two assets classes are highly popular, their risk transmission role suggests that they should be complemented with strong hedging assets. Besides that, given the ineffective hedging of AI tokens with other assets, investors should closely monitor their exposure to AI-related cryptos and adjust their allocations accordingly. MCP outperforms MVP and MCoP, making it the most suitable strategy for investors, especially those interested in AI tokens. Second, given the risk-spreading nature of AI ETFs and clean energy assets, regulators should closely monitor their volatility and systemic impact. Implementing safeguards to prevent excessive market fluctuations is essential. Additionally, policymakers should support financial instruments that facilitate the integration of these assets into broader portfolios, preventing them from becoming hidden sources of risk.

### v2

@@ 6.2 Multivariate portfolio analysis
Table 3 show the results for MVP, MCP, MCoP, $\text{MCoP}^{C}$, and $\text{MCoP}^{L}$. It is worth emphasizing that MCoP, $\text{MCoP}^{C}$, and $\text{MCoP}^{L}$ are derived from total, contemporaneous, and lagged connectedness, respectively. We first discuss the performance of AI ETFs in the MVP strategy. The relatively high HE values (IRBO: 0.905, ROBT: 0.890) indicate that multivariate portfolios are effective in reducing the risk of AI ETFs. Even though AI ETFs are the primary transmitters of shocks in the return transmission analysis, their weights remain relatively low (IRBO: 2.4%, ROBT: 3.8%). This suggests that their role in the MVP portfolio is limited. In the MCP, the allocation to AI ETFs increases significantly (IRBO: 6.0%, ROBT: 12.5%), indicating an enhanced contribution to the portfolio. However, their HE values turn negative (IRBO: $-2.215$, ROBT: $-2.728$), indicating that this portfolio fails to effectively mitigate volatility in AI ETFs and may even increase risk. In the MCoP method, the allocation to AI ETFs further increases (IRBO: 10.5%, ROBT: 11.5%), reinforcing their role within the portfolio. Nevertheless, their HE values continue to decline (IRBO: $-3.375$, ROBT: $-4.073$). Furthermore, the MCoP , $\text{MCoP}^{C}$, and $\text{MCoP}^{L}$ allocation methods assign very close weights to AI and green assets. Overall, despite Figs. 2, 4 and 6 indicating that AI ETFs are major risk spillovers, their weights in these portfolios remain relatively limited.
@@
Next, we analyze the role of AI Tokens across different portfolio strategies. In the MVP, AI tokens (FET, FIL, ICP, NEAR, RENDER) have minimal weights, each below 1%, limiting their overall impact on the portfolio. Furthermore, their HE values exceed 0.99, suggesting that this investment strategy effectively minimizes their risk. In the MCP portfolio, despite a decrease in HE values (ranging between 0.783 and 0.854), the weights of AI Tokens increase significantly (e.g., RENDER: 12.6%, FET: 11.6%), indicating their enhanced role within the portfolio. In the MCoP, the weights of AI Tokens further increase, reflecting their rising importance in this investment strategy. Their HE values remain above 0.7 (ranging between 0.705 and 0.802), suggesting that this strategy contribute to volatility reduction of AI tokens, though to a lesser extent than in the MVP and MCP. Moreover, portfolios constructed using $\text{MCoP}^{C}$ and $\text{MCoP}^{L}$ also exhibit a similar pattern. Overall, AI Tokens exhibit higher allocation ratios for both MCP and MCoPs, whereas their weights in the MVP remains low due to higher volatility, limiting their overall influence.
@@
Finally, we examine the performance of green assets in the portfolios. Green assets remain the core holdings in the MVP, with the highest allocation. SPGB holds an overwhelming weight of 91.2%, while SPGTCED accounts for 0.7%, likely reflecting their respective roles as a risk receiver and a risk contributor. Accordingly, SPGTCED has an HE value of 0.884, demonstrating a significantly stronger risk reduction effect compared to SPGB, which has a negative HE value of --0.05. From the MCP procedure, the weight of green assets declines significantly (SPGB: 19.7%, SPGTCED: 13.2%, total: 32.9%), and their HE values continue to decline, indicating a weakening risk reduction in this portfolio. In the MCoP, the allocation to green assets decreases further (SPGB: 11.2%, SPGTCED: 11.0%, total: 22.2%), with HE values continuing to drop, suggesting a further reduction in their role. Despite Figures 2, 4 and 6 demonstrating that clean energy transmits shocks to other markets, its weight in the portfolio is not dominant. Again, the results from $\text{MCoP}^{C}$ and $\text{MCoP}^{L}$ are very close to those of MCoP. This finding aligns with previous literature (Tiwari et al., 2022), which also found that while clean energy is a risk contributor, its allocation in investment portfolios is not necessarily the highest, whereas green bonds tend to account for a larger proportion of portfolio allocation.
Interestingly, in the case of the MVP, asset allocations, except for AI tokens, seemly correspond to their roles. Risk receivers have relatively higher weights, while risk transmitters have lower weights. The results of MCP and MCoPs are quite similar, possibly because both focus on reducing correlations across assets. All the portfolios reduce the risk of AI tokens but increase volatility of the green bond.
6.3 Portfolio performance
We illustrate the cumulative portfolio returns in Fig. 8. This figure presents the cumulative returns of portfolio strategies. MVP remains relatively stable with minimal fluctuations, indicating a low-risk but low-return profile, while MCP and MCoPs show significant growth, especially from early 2023 onward. In the early period, all portfolios experience negative returns, but MCP and MCoPs recover strongly, whereas MVP remains around the zero level. MCP and MCoP follow similar trends, with MCoP generally achieving slightly higher peaks, suggesting better performance in specific market conditions. Other studies on investment strategy comparisons also documents negative cumulative portfolio returns (Rubbaniy et al., 2024; Lang et al., 2023).
Figure 8: Cumulative returns of different portfolios.
@@ 6.3 Portfolio performance
SPGB $0.197$ $0.034$ $0.139$ $0.261$ $-34.611$ $0.000$
SPGTCED $0.132$ $0.034$ $0.061$ $0.176$ $-2.927$ $0.000$
MCoP IRBO $0.105$ $0.112$ $0.000$ $0.250$ $-3.375$ $0.000$
ROBT $0.115$ $0.120$ $0.000$ $0.250$ $-4.073$ $0.000$
@@
SPGB $0.112$ $0.118$ $0.000$ $0.250$ $-47.457$ $0.000$
SPGTCED $0.110$ $0.113$ $0.000$ $0.250$ $-4.343$ $0.000$
MCoPC IRBO $0.116$ $0.115$ $0.000$ $0.250$ $-3.426$ $0.000$
ROBT $0.105$ $0.116$ $0.000$ $0.250$ $-4.132$ $0.000$
FET $0.112$ $0.119$ $0.000$ $0.250$ $0.800$ $0.000$
FIL $0.104$ $0.111$ $0.000$ $0.250$ $0.714$ $0.000$
ICP $0.113$ $0.115$ $0.000$ $0.250$ $0.701$ $0.000$
NEAR $0.109$ $0.119$ $0.000$ $0.250$ $0.704$ $0.000$
RENDER $0.113$ $0.111$ $0.000$ $0.250$ $0.795$ $0.000$
SPGB $0.115$ $0.116$ $0.000$ $0.250$ $-48.025$ $0.000$
SPGTCED $0.113$ $0.120$ $0.000$ $0.250$ $-4.406$ $0.000$
MCoPL IRBO $0.105$ $0.113$ $0.000$ $0.250$ $-3.753$ $0.000$
ROBT $0.121$ $0.124$ $0.000$ $0.333$ $-4.512$ $0.000$
FET $0.109$ $0.114$ $0.000$ $0.250$ $0.785$ $0.000$
FIL $0.106$ $0.110$ $0.000$ $0.250$ $0.693$ $0.000$
ICP $0.106$ $0.119$ $0.000$ $0.250$ $0.679$ $0.000$
NEAR $0.118$ $0.113$ $0.000$ $0.250$ $0.682$ $0.000$
RENDER $0.112$ $0.114$ $0.000$ $0.250$ $0.780$ $0.000$
SPGB $0.116$ $0.111$ $0.000$ $0.250$ $-51.654$ $0.000$
SPGTCED $0.108$ $0.116$ $0.000$ $0.250$ $-4.806$ $0.000$
Finally, the Sharpe ratios of the three portfolios are shown in Table 4. The MCP emerges as the most optimal strategy, as it achieves the highest return and Sharpe ratios across all evaluated metrics. While the MVP effectively minimizes volatility, its negative return and Sharpe ratios render it an unattractive investment option. $\text{MCoP}^{C}$, and $\text{MCoP}^{L}$ have negative return and Sharpe ratios. The MCoP offers a moderate risk-return profile, yet it fails to surpass MCP. Therefore, for investors seeking an optimal balance between risk and return, the MCP portfolio represents the most advantageous selection. This finding is similar with Ziadat et al. (2024), who also observe a negative Sharpe ratio.
@@
MVP MCP MCoP MCoPC MCoPL
Return $-0.0165$ $0.2664$ $0.2438$ $-0.0335$ $-0.0932$
StdDev $0.0835$ $0.4860$ $0.5669$ $0.5702$ $0.5910$
Sharpe Ratio (StdDev) $-0.1978$ $0.5482$ $0.4300$ $-0.0588$ $-0.1578$
Sharpe Ratio (VaR) $-1.9865$ $5.8478$ $4.8230$ $-0.6228$ $-1.7048$
Sharpe Ratio (CVaR) $-1.4889$ $4.2027$ $3.8360$ $-0.4194$ $-1.2778$
@@ 7 Conclusion
This study employs the $R^{2}$ decomposition to examine the risk spillover dynamics among AI ETFs, AI tokens, and green assets over the period from May 25, 2021, to January 31, 2025. Our findings reveal that contemporaneous spillovers dominate the total connectedness index, while lagged spillovers remain relatively stable. Furthermore, AI ETFs and clean energy emerge as risk transmitters, whereas AI tokens and green bonds primarily act as risk receivers. In addition, AI tokens are challenging to hedge and offer little benefit in hedging other assets when considering bilateral hedge ratios. However, clean energy and AI ETFs exhibit stronger hedging effectiveness regardless of the long or short position taken. Additionally, we analyze the performance of the following portfolio strategies: MVP, MCP, and MCoPs. The results show that while all the approaches reduce volatility of AI tokens significantly, they increase the risk associated with green bonds. However, only the MVP decreases the risk of AI ETFs and clean energy, while the other strategies actually increase their risk. Finally, MCP exhibits superior performance than the other portfolios. In this investment strategy, green bonds have the largest allocation, while IRBO has the smallest.
The nature of these assets as financial instruments might be helpful to better understand their risk transmission. Although AI tokens are thematically linked to artificial intelligence, they fundamentally belong to the category of crypto assets, which are often speculative, volatile, and sensitive to external shocks (Kukacka and Kristoufek, 2023; Osman et al., 2024). Their highly synchronized price movements suggest internal spillovers and clustering effects.
On the contrary, AI ETFs represent firms with actual research and development capabilities in AI. The pricing of these assets is more directly driven by market expectations regarding AI advancement and earnings potential. Spillovers from AI ETFs to green assets may partly reflect market perceptions of AI’s enabling role in advancing environmental sustainability, as suggested by existing literature (Vinuesa et al., 2020; Entezari et al., 2023; Nahar, 2024).
Green bonds, as relatively stable and policy-backed debt instruments, act as risk receivers within the financial system, consistent with findings of Tiwari et al. (2022). In contrast, clean energy equities are characterized by technological orientation and strong growth potential. Zeng et al. (2024) and Tiwari et al. (2022) also indicate that they may serve as sources of risk spillovers in the system. Such spillovers may, at least in part, reflect the broader influence of clean energy technologies. Therefore, future research could empirically examine how asset-specific characteristics and underlying structures influence the risk spillover across different types of assets.
Our work has the following implications: First, given the risk-spreading nature of AI ETFs and clean energy assets, regulators should closely monitor their volatility and systemic impact. Implementing safeguards to prevent excessive market fluctuations is important. Additionally, policymakers should pay close attention to the interconnectedness between technology-related financial instruments and green assets, particularly within the frameworks of sustainable finance and climate risk management. In this context, enhanced cross-market risk monitoring and response mechanisms are essential. Moreover, aligning technology policy with green finance policy may be necessary to effectively address the systemic implications of such interlinkages.
@@
Second, investors should consider the high spillover risks associated with AI ETFs and clean energy. Although the two assets classes are highly popular, their risk transmission role suggests that they should be complemented with strong hedging assets. Besides that, given the ineffective hedging of AI tokens with other assets, investors should closely monitor their exposure to AI-related cryptos and adjust their allocations accordingly. MCP outperforms MVP and MCoPs, making it the most suitable strategy for investors, especially those interested in AI tokens. Investors should also be mindful of potential herding behavior among AI tokens, as their highly synchronized dynamics and strong interconnections may pose hidden systemic risks.